a thin glass prism of refractive index 1.5 produces a deviation of 4 of a ray incident at a small angle . what will be he deviation of the same incident ray by the same prism if the it is immersed in water of refractive index 4/3

Grade:11

1 Answers

Hiral
5 years ago
Let the angle of prism be A 
Deviation =(u-1)A
4°=(1.5-1)A
So we get A=8° now when we dip in water so refractive index of water w.r.t glass ug/uw=3/2/4/3=9/8
Deviation =(9/8-1)8=1°
Hence answer=1°